Let f(x)= 5x-1 and g(x)= x^2-5
Find (fog)(-2)

Answer by stanbon(75887)

Let f(x)= 5x-1 and g(x)= x^2-5
Find (fog)(-2) = f[g(-2)] = f[(-2)^2-5] = f[-1] = 5(-1)-1 = -6
--------------
Cheers,
Stan H.
------------